OpenStudy (anonymous):

PLEASE HELP ME How many solutions are there to the following system of equations? 2x - y = 2 -x + 5y = 3 A. 0 B. 1 C. 2 D. infinitely

OpenStudy (anonymous):

@Ellia201 @emc726 @AaronAndyson

OpenStudy (anonymous):

hold up

OpenStudy (horsegirl27):

What math is this?

OpenStudy (anonymous):

no i think its two only

OpenStudy (anonymous):

13/9 and 8/9

Nnesha (nnesha):

use elimination graphing matrices or substitution which one is easy for u ???

OpenStudy (anonymous):

so c

OpenStudy (anonymous):

its Algebra 1

OpenStudy (horsegirl27):

Ok I take algebra 1 and I recognize this

OpenStudy (horsegirl27):

I think I know how to do it one second...

OpenStudy (anonymous):

2x-y=2 sove for y ..... y=2x-2 -x+5y=3 -x+2x-2=3 ===> x=5 y=2*5-2=8

OpenStudy (anonymous):

i swear to god if my answer is right then uall are wasting ur time

OpenStudy (mrnood):

These equations are for 2 straight lines. If two lines are parallel they do not intersect, and therefore there is no solution to the 'system' If th ey are not parallel then the can only intersect at ONE point. You can tell if they are parallel - convert them both to the form y=mx + c m is th gradient - if they have the same gradient then they ARE parallel

OpenStudy (horsegirl27):

2 equal signs?

OpenStudy (mrnood):

@MSM6205 It is IMPOSSIBLE for two straight lines ot intersect at TWO places. Your answer CANNOT be correct. Since oyu gave NO explanation we will never know where you went wrong.

OpenStudy (anonymous):

like what @Nnesha

Nnesha (nnesha):

look at ur answer choices and question again seem like** there is something wrong

OpenStudy (anonymous):

@MrNood first of all usually my answers are not wrong second of all u can check this up online on mathway.com

OpenStudy (mrnood):

@MSM6205 'Usually' does not matter two straight lines cannot cross in two places

Nnesha (nnesha):

bec both lines are not parallel after solving any method either elimination substitution or graphing the answer i got is not up there

OpenStudy (mrnood):

@MSM6205 If oyu care to show how you think oyu got 2 solutions then I will show you wher you went wrong

OpenStudy (mrnood):